The conjugate base of NH3 is NH3+. yes or no?

Grade:12

4 Answers

Ashwin Muralidharan IIT Madras
12 years ago

Hi Harish,

 

NH4+ is the conjugate Acid of NH3...

It can't be NH3+.

 

Best Regards,

Ashwin (IIT Madras).

basit ali
12 years ago

YES, because when a molucule gain proton it is caled conjugate base.

  given reaction folows this equation:

    NH3 →H- + NH3+   

Godfrey Classic Prince
12 years ago

Dear Harish Kumar.M,

Ammonia is amphoteric, meaning that it can behave as either an acid or a base. It can accept a proton and therefore acts as a base. It can donate a proton and act as an acid.

NH3 + H2O <==> H3O+ + NH2-

NH2- is therefore the conjugate base of ammonia.

 

All the Very Best & Good Luck to you ...

Hope this helped you immensely...

Regards,

AskIITians Expert,

Godfrey Classic Prince

IIT-M

Please approve my answer if you liked it by clicking on "Yes" given below...!!!Smile

Arka
6 years ago
No not possible NH3+ because NH4+ is the conjugate acid of NH3NH3(aq)+H2O(l)=NH4+(aq)+OH-(aq)Therefore,base+acid=conjugate+conjugate acid baseHence,base adds proton to conjugate acid.And Acid looses proton to form conjugate base